Phosphorus pentachloride is formed when 17.2 g of chlorine gas react with 23.2 g of solid phosphorus (P2). Determined the reactant that is in excess


1
Expert's answer
2022-02-19T06:57:55-0500

Here is the equation:

P2 + 5Cl2 --> 2PCl5


"n(P_2)=23.2\ g\times\frac{1\ mol}{61.95\ g}=0.374\ mol"

"n(Cl_2)=17.2\ g\times\frac{1\ mol}{70.91\ g}=0.243\ mol"


According to the equation, P2 and Cl2 react in 1 : 5 molar ratio. The actual ratio is 0.374 : 0.243, which indicates the significant excess of P2.


Answer: P2